Calculate the molarity of 2.60 g of sodium chloride in 40.0 ml of solution 


1
Expert's answer
2021-01-11T04:44:48-0500

M(NaCl) = 23+35 = 58 g/mol

n(NaCl) "= \frac{2.6}{58} = 0.0448 \;mol"

Proportion:

0.0448 mol – 40 mL

x – 1000 mL

"x = \frac{0.045 \times 1000}{40} = 1.12 \;M"

Answer: 1.12 M